OpenStudy (anonymous):

HELP PLZ What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1 A. -69x + 10y = 51 B. 69x + 10y = -51 C. -69x + 10y = -51 D. 69x + 10y = 51

OpenStudy (anonymous):

move -6.9x to the other side and multiply the whole thing by ten, you should see the answer :)

OpenStudy (anonymous):

What all the answers have in common is that they multiplied everything by 10. So let's do that: 10y = -69x + 51 Then we want to move the x and y terms to one side. We do this by adding 69x to both sides: 10y + 69x = -69x +69x + 51 (The two 69x terms on the right cancel out, something - itself is naturally 0) 10y + 69x = 51 Switch the x and y terms around: 69x + 10y = 51 The answer is D.

OpenStudy (anonymous):

great explaining @IAmSinged! thank you both :) mind helping with a couple more?

OpenStudy (anonymous):

Go ahead and post them.

OpenStudy (anonymous):

okay ty :)

OpenStudy (anonymous):

@alias @IAmSinged

OpenStudy (anonymous):

which one pass the vertical line test? Do you know what the test is?

OpenStudy (anonymous):

if you don't, just run a ruler (or anything straight) vertically through the graph, and if it hits two or more points, then it fail

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Confirming B fails.

OpenStudy (anonymous):

:S

OpenStudy (anonymous):

hmm D @alias

OpenStudy (anonymous):

yup, last one is right :)

OpenStudy (anonymous):

yayy! :D

OpenStudy (anonymous):

if you notice, for b, when you move in the line, you have infinite number of points, so that violate the vertical line test